OpenStudy (anonymous):

Find the equation for a line which passes through the points (–1, 2) and (–3, 8). Write the equation in slope-intercept form

OpenStudy (anonymous):

y = mx +b m = (y2-y1)/(x2-x1) y2 = 8 y1 = 2 x2 = -3 x1 = -1 plug in the slope and one of the ordered pairs (x,y) into the equation to solve for b.

OpenStudy (aravindg):

Hi, I can see that you are a new member here at OpenStudy so, I would first like to say Welcome to OpenStudy, I would like to point you to the chat pods these chat pods are where you can make new friends and talk to new people just like you, I would also like to emphasize our "NON-CHEATING POLICY" we ask here at OpenStudy that you do NOT post exam/Test questions if you are caught doing this we will notify your current school, I would also like to ask you to check out the OpenStudy "Code of Conduct" http://openstudy.com/code-of-conduct Please read this carefully and thoroughly. Welcome to the OpenStudy Community! Ambassador

OpenStudy (anonymous):

\[y-8/x-(-3)=8-2/-3(-1)\]